OpenStudy (anonymous):

The air in a container has a constant temperature of 43.7° F. The air is then cooled at a rate, measured in degrees Fahrenheit, represented by the function R(x) = 1.4x - 1, where x is the time in hours. Which function best describes A(x), the temperature of the air over time? A(x) = 44.7 - 1.4x A(x) = 42.7 + 1.4x A(x) = 43.7(1.4x) - 43.7 A(x) = 1.4x - 44.7

OpenStudy (perl):

Welcome to Openstudy

OpenStudy (anonymous):

Thank you :)

OpenStudy (perl):

you want to find A(x)

OpenStudy (anonymous):

I have no clue of how to write the function

OpenStudy (perl):

i can solve this using calculus, but are you learning calculus?

OpenStudy (anonymous):

no

OpenStudy (perl):

may I ask where did you get this question.

OpenStudy (anonymous):

study island

OpenStudy (perl):

make sure I am copying this correctly the rate at which it is cooling is R(x) = 1.4*x - 1

OpenStudy (anonymous):

yes

OpenStudy (perl):

I am asking all these questions because the question seems odd. None of the choices match what I think is the correct answer

OpenStudy (anonymous):

that is what is throwing me off...

OpenStudy (perl):

This is one way to solve this, but it doesn't match your choices. Using calculus we can integrate the rate R(x) to get the temperature A(x) = integral (1.4x - 1) dx A(x) = 1.4 x^2/2 -x + C A(0) = 43.7 A(x) = 1.4 x^2/2 - x + 43.7 @ganeshie8 maybe you could look this over

OpenStudy (anonymous):

that would make sense but i am limited to their choices.

OpenStudy (perl):

If i had to guess i would pick a)

OpenStudy (perl):

since a) is the only option with a negative slope or rate (since it is cooling)

OpenStudy (perl):

is it possible for you to upload the question, so I can see the original

OpenStudy (perl):

if you can take a screenshot

OpenStudy (anonymous):

the correct answer turned out to be B

OpenStudy (anonymous):

Lynn works at a retail department store where she has a coupon to receive $10.00 off her next purchase. She also receives a 31% employee discount. If p represents the amount of her purchase, which function applies the coupon and discount so that Lynn pays the lowest amount? D(p) = 0.69p - 10 D(p) = 0.31(p - 10) D(p) = 0.31p - 10 D(p) = 0.69(p - 10) can you make any sense out of this one?

OpenStudy (perl):

b) is wrong. they did A(x) = A(0) + R(x)

OpenStudy (perl):

which doesn't make sense.

OpenStudy (perl):

yes we can do the coupon problem :)

OpenStudy (anonymous):

im starting to think study island is rigged against me...

OpenStudy (anonymous):

i would think the first option

OpenStudy (perl):

There are two ways she can pay for the item Option 1: she can first apply the ten dollar coupon, then the 31% employee discount: (p - 10)* ( 1 - .31) = (p - 10) * .69 = Option 2: first apply the 31% employee discount, then ten dollar coupon: p( 1 - .31) - 10 = .69*p - 10 Let's plug in p = 100 to see which comes out less. option 1: (100 - 10) * .69 = 62.1 option 2: .69 * 100 - 10 = 59

OpenStudy (perl):

so option 2 is less D(p) = .69p - 10

OpenStudy (perl):

which corresponds to a) in your multiple choices

OpenStudy (anonymous):

thank you :)

OpenStudy (perl):

is it correct ?

OpenStudy (anonymous):

how about this one? Currently, there are 93 bald eagles in Forest National Park. Bald eagles are an endangered species, and the population of bald eagles is decreasing by 7% every year. Which of the following functions represents the number of bald eagles in Forest National Park in t years? B(t) = 93(0.93)t B(t) = 93(1.07)t B(t) = 93(-0.93)t B(t) = 86.49t

OpenStudy (perl):

oh i see option 1 is equivalent to (p - 10)* ( 1 - .31) = (p - 10) * .69 = = .69*p - 10*.69 = .69p - 6.9 option 2 is .69p - 10 so you can see option 2 is lower

OpenStudy (perl):

we can use the function y = 93 ( 1- r / 100)^t , where r is given as a percent , t is in years, y represents the remaining eagles